### v1

@@ Abstract
Many studies have examined social determinants of health (SDoH) factors independently, overlooking their interconnected and intersectional nature. Our study takes a multifactorial approach to construct a neighborhood-level measure of SDoH and explores how neighborhood residency impacts care received by endometrial cancer patients in Massachusetts. We used a Bayesian multivariate Bernoulli mixture model to create and characterize neighborhood SDoH (NSDoH) profiles using the 2015–2019 American Community Survey at the census tract level (n=1478), incorporating 18 variables across four domains: housing conditions and resources, economic security, educational attainment, and social and community context. We linked these profiles to Massachusetts Cancer Registry data to estimate the odds of receiving optimal care for endometrial cancer using Bayesian multivariate logistic regression. The model identified eight NSDoH profiles. Profiles 1 and 2 accounted for 27% and 25% of census tracts, respectively. Profile 1 featured neighborhoods with high homeownership, above-median incomes, and high education, while Profile 2 showed higher probabilities of limited English proficiency, renters, lower education, and working-class jobs. After adjusting for sociodemographic and clinical characteristics, we found no statistically significant association between NSDoH profiles and receipt of optimal care. However, compared to patients in NSDoH Profile 1, those in Profile 2 had lower odds of receiving optimal care [OR = 0.77, 95% CI (0.56, 1.07)]. Our results demonstrate the interconnected and multidimensional nature of NSDoH, underscoring the importance of modeling them accordingly. This study also highlights the need for targeted interventions at the neighborhood level to address underlying drivers of health disparities, ensure equitable healthcare delivery, and foster better outcomes for all patients.
@@ 1 Background
Social determinants of health (SDoH) are increasingly recognized as critical factors that shape healthcare outcomes by influencing access to care, treatment decisions, and overall patient experiences[1, 2, 3]. These determinants, including socioeconomic status, experiences of racism or discrimination, and neighborhood conditions, can introduce bias in the care that individuals receive, potentially leading to disparities in treatment and outcomes[4]. Endometrial cancer (EC), one of the most commonly occurring female cancers, serves as an example of how SDoH can influence care[5, 6]. Despite the existence of evidence-based treatment guidelines, disparities across the EC care continuum persist, particularly among ethnically and socioeconomically disadvantaged groups[7, 8, 9, 10, 11, 6, 5, 12, 13]. A comprehensive analysis of SDoH that transcends race and ethnicity would provide researchers and policymakers with a deeper understanding of how the social environment can influence a patient’s ability to engage with the healthcare system [11, 12].
@@
In settings where individual-level SDoH measures are not readily available, researchers often rely on non-standardized composite scores of proxy variables. These variables correspond to different domains of SDoH, such as income, education, and employment[14, 15]. At the geographic level, SDoH has been operationalized using area-based measures derived from publicly available data sources such as the United States (US) Census and the American Community Survey (ACS). Two commonly used measures are the Yost Index and the Area Deprivation Index (ADI)[16, 17, 18]. The Yost Index, employed in national cancer surveillance, currently operates under the assumption that SDoH is one-dimensional. It inputs seven area-level variables (median household income, education, proportion of households below the 150% poverty line, median house value, median gross rent, unemployment rate, and working-class occupation). These variables are summarized at the census tract or block level through a Principal Component Analysis (PCA). Here, the primary component, which accounts for the most variance among the variables, is selected as the socioeconomic status (SES) index.[16, 19, 20]. The component loadings are used to calculate a composite SES score for each areal unit, and these are further ranked based on their SES scores and divided into quintiles, with quintile 1 representing the lowest SES and quintile 5 the highest.
@@
The ADI, originally developed in 2002 to study general mortality disparities and hospitalizations in the US, expands the set of census-derived SDoH variables to 17. Previously, it operated under the same one-dimension assumption as Yost, but now incorporates a three-factor model to accommodate multi-dimensionality [21, 20, 18, 17]. Highly correlated variables are first clustered together and then defined along the following three domains: financial strength, economic hardship and inequality, and educational attainment [18]. As in the Yost Index, areal units are indexed based on the loadings of each of these three-factor domains, where higher percentile ranks indicate greater socioeconomic disadvantage. Boscoe et al. compared these two indices and found that despite being derived from the same data and heavily influenced by financial domains, they differ significantly in terms of transparency for reproducibility, distribution, and sensitivity to missing data[20]. Methodologically, these approaches share similarities in their use of principal components or factor analysis to reduce the number of highly correlated SDoH variables, but the domains derived are treated independently from one another and fall short of acknowledging the ways in which they intersect.
@@
Recently, Kolak et al. introduced a new approach that further expands the list of census-derived SDoH variables for PCA, including variables like limited English proficiency [3]. These variables were reduced to four domains: socioeconomic advantage index, limited mobility index, urban core opportunity index, and mixed immigrant cohesion and accessibility index. Additionally, they clustered census tracts by characterizing them into profiles based on similar scoring across the four SDoH components using K-means clustering. While this approach does allow for clustering based on shared responses to the different domains, the clustering is based on a single summary score. Overlooking how each component-specific score contributes to the summary score ignores the heterogeneous ways in which area deprivation can manifest. Two neighborhoods could potentially be clustered together but experience different patterns of resource allocation, which could ultimately influence how those residents interact and are able to access healthcare services. New methodological approaches are needed to accommodate the interconnectedness and multifactorial structure of SDoH, inclusive of factors pertinent to vulnerable populations that are at greater risk of health disparities[20, 3, 14, 22].
@@
Model-based clustering methods, specifically finite mixture models (FMM), are able to identify underlying patterns in heterogeneous populations from a wide set of interrelated variables. Lekkas et al. recently highlighted the methodological advantages of FMMs in neighborhood health research through the common application of Latent Class Analysis (LCA) and Latent Profile Analysis (LPA). These models are able to characterize neighborhood profiles based on a set of multivariate categorical or continuous data and examine their association with various health indicators [23]. Although LCA or LPA are well-suited to handle the multi-faceted nature of neighborhoods, typically, the number of latent patterns/profiles is unknown, and multiple model fitting and testing is required to determine the optimal number of classes. However, under a Bayesian framework, we are able to rely on data-driven algorithms to identify an interpretable and reproducible set of neighborhood SDoH (NSDoH) profiles with the added flexibility of incorporating prior information into parameter estimation [24]. These models have found great utility in nutrition, environmental health, and population genomics [25, 26, 27, 28, 29, 30, 31]. Yet, none have been applied to area-level exposures of SDoH.
@@
In this study, we aim to generate and characterize new NSDoH profiles to better understand their impact on EC care in Massachusetts. Given binarized aggregate proportions from the ACS, we will implement a Bayesian multivariate Bernoulli mixture model(MBMM) to cluster neighborhoods that share similar SDoH characteristics. Neighborhoods are established based on census tract identifiers, as defined by the U.S. Census Bureau from 2015 to 2019. Once identified, we will use these profiles with patient-level data from the Massachusetts Cancer Registry to determine if the receipt of optimal care for EC differs by a patient’s residential neighborhood.
@@
We organize the remaining sections of this paper as follows. In section 2, we describe our methodological approach. In section 3, we share our results of the MBMM and examine its association with EC treatment. Finally, in sections 4 and 5, we provide a discussion and concluding remarks.
@@ 2.1 Data Source
We used census tract (neighborhood) level aggregate data from the 2015-2019 five-year American Community Survey (ACS) estimates for the state of Massachusetts (MA). The data contained a total of 1478 census tracts that included approximately 6.8 million persons. Based on previous research, we selected eighteen(18) variables as indicators of NSDoH [20, 3]. Given that the selected NSDoH variables were aggregate proportions and the Bernoulli distribution models binary data, we dichotomized each variable based on whether that feature fell above $(x_{ij}=1)$ or below $(x_{ij}=0)$ the state median value. Table 1 details the median threshold determined for each of the SDoH features. Neighborhoods above that specific threshold were classified as high exposure. Neighborhoods below that specific threshold were classified as low exposure. Data was extracted from respective ACS tables via the tidycensus package implemented in the R software environment [32].
[…]
@@ 2.2 Multivariate Bernoulli Mixture Model (MBMM)
Neighborhood social determinants of health (NSDoH) profiles were identified using a fully Bayesian estimation of a multivariate Bernoulli mixture model (MBMM). This model clusters a population based on shared responses to a set of binary exposures. Let $i\in\{1,...,n\}$ index an individual census tract, where $n$ is the total number of census tracts or neighborhoods in Massachusetts. Let $\mathbf{x_{i}}=\{x_{i,1},..,x_{i,p}\}$ denote a vector of observed binary indicators, where $x_{i,j}=1$ denotes a high exposure (above state median threshold) to social determinant $j\in\{1,..,p\}$ in census tract $i$. Let $K$ denote the number of NSDoH profiles in the population. In practice, this number is typically not known a priori. To determine an appropriate number of profiles in the model, we overfit the model with $K_{max}$ clusters, which over-exceeds the true number of clusters and allows the observed data to drive the number of nonempty clusters estimated at the end of the Bayesian sampling algorithm [44]. The observed likelihood for the set of binary data $\mathbf{X}=\{0,1\}^{n\times p}$ under the MBMM is given by
@@ 2.2.1 Posterior Computation
Model parameters are estimated via a Metropolis-coupled Markov chain Monte Carlo ($MC^{3}$) algorithm. An allocation sampler is used to determine the most probable number of NSDoH profiles. This is based on the maximum a posteriori number of nonempty clusters from the posterior distribution,$K_{map}$. Once $K_{map}$ is determined, we calculate the posterior mean of each of the model parameters. These posterior estimates are then used to describe each NSDoH profile $k$, based on the posterior distribution of $\boldsymbol{\theta_{\cdot|k}}=(\theta_{1|k},\ldots,\theta_{p|k})$, and assigning each MA census tract to the profile with the greatest posterior probability of membership. Further details on prior selection, model fit testing, and other cluster postprocessing procedures are provided in Supplemental Materials A.
@@ 2.2.2 Regression Analysis
Patient-level data was obtained from the Massachusetts Cancer Registry (MCR) of the Massachusetts Department of Public Health (MDPH). A total of 2412 records were collected for females, aged 18 years or older and diagnosed with endometrial cancer between 2015 and 2017. Cases were identified using the International Classification of Diseases for Oncology, third edition (ICD-O3) primary site (C54.1) and morphology codes for endometrial carcinoma [45]. The main outcome variable was adherence to National Comprehensive Cancer Network (NCCN) guidelines for the first course of treatment while accounting for guideline changes during the study period. These guidelines are based on tumor, grade, and stage, incorporating a combination of surgery, chemotherapy, and radiation as necessary. Adherence to NCCN treatment guidelines was treated as a dichotomous variable where we defined optimal care as the patient receiving therapy in accordance with NCCN guidelines. The main exposure was NSDoH profiles derived from the MBMM in 2.2 and assigned to each patient’s residential census tract at EC diagnosis. Associations were examined by adjusted odds ratio using a Bayesian logistic regression, adjusting for patient-level clinical and sociodemographic characteristics at cancer diagnosis (year of diagnosis, age, insurance status, and initial point of care facility type). Additionally, we conducted a supplementary analysis to examine the association between NSDoH profiles and the type of initial care facility (academic medical center versus other facilities). More details about the guidelines and regression analyses are provided in the Supplemental Materials Section B.
@@ 3.1 NSDoH Profiles Results
Table 1 shows the NSDoH variables that were selected from the 2015-2019 ACS survey and clustered to characterize census tracts in Massachusetts. MA census tracts were mostly composed of owner-occupied households and high proportions of $\geq$ HS and $\geq$ Bachelor’s degree educational attainment. On average, across all census tracts, about 13% of households received SNAP benefits in the past 12 months, and 52% of the population aged 16 or older had non-white collar occupations.
@@
The model identified eight NSDoH profiles. Figure 1 illustrates the estimated posterior probability of high exposure to each of the 18 NSDoH variables grouped by domain, given assignment to that specific NSDoH profile. Profile 1 represented 27% of MA census tracts. This profile is characterized by neighborhoods that are more likely to have owner-occupied housing (99.7%), income above the state median (95.8%), and higher than high school education (95.6%). It also had a lower probability of exposure to working-class employment (4.7%), participation in federal assistance programs (6.6%), and most social and community context indicators. Profile 2 represented the second largest cluster, containing 25% MA census tracts. This cluster had the highest probability of exposure to households with limited English proficiency (96.4%), female head of households (90.8%), and household crowding (84.1%). High probabilities of exposure were also identified for neighborhoods with a lack of complete plumbing (39.5%), renter-occupied housing (99.7%), no vehicle access (94.6%), working class (98.1%), and all three ethnic minority groups, especially the proportion of non-Hispanic Black (81%) and Hispanic/Latino residents (94.5%). This profile also had the lowest education levels, with almost all reporting less than HS education (98.4%), compared to all other NSDoH profiles. Overall, this profile exhibited characteristics of disadvantage and deprivation across all thematic domains. Profile 3 favored neighborhoods with high probabilities of exposure to housing conditions and resources variables (renter-occupied, no vehicle, household crowding), high proportions of ethnic minorities (51-83%), and educational attainment (HS and Bachelor’s degree education). Profile 4 shared similarities with NSDoH profile 1 along the housing conditions and resources domains but differed in economic security (more working class) and education (less Bachelor’s degree education). It also had the lowest probabilities of exposure for social and community context indicators compared to all other profiles. Profile 5 favored neighborhoods with high probabilities of exposure to owner-occupied housing, less than HS education, and working class, moderate probability of ethnic minorities (48-68%), and low probabilities for lack of complete plumbing (21.4%) and no vehicle (30.2%). Profile 6 favored neighborhoods with high probabilities of exposure to higher proportions of ethnic minority groups (85-95%), residents with limited English proficiency (94.5%), renter-occupied housing (98.9%), no vehicle in the household (96.5%), and SNAP benefits (90.8%). Although similar to Profile 2, a notable observation is a divide in educational attainment, particularly between those with less than HS education (85.8%) and those with at least a Bachelor’s degree education (96%). Profile 7 shared similarities with NSDoH profile 5, but differed in education and social and community context variables. Profile 7 had higher probabilities of exposure to those with high school and bachelor’s degree education and lower probabilities of exposure to the proportion of ethnic minority residents (29%-33%) and limited English proficiency (17%). Lastly, Profile 8 was characterized by neighborhoods with a higher probability of exposure to residents with less than HS education (47.1%), households living below the poverty line (84.4%), participation in federal assistance programs (82.7%), renter-occupied housing (98.7%), and working-class employment (80.8%), and a lower probability of exposure to households with income above the state median (10.2%). Neighborhoods in this profile had a 60% probability of being exposed to residents with a female head of household and close to 50% estimated probability for an above state median proportion of Hispanic/Latino or non-Hispanic Black residents.

### v2

@@ Abstract
Many studies have examined social determinants of health (SDoH) independently, overlooking their interconnected nature. Our study uses a multidimensional approach to construct a neighborhood-level measure that explores how multiple SDoH jointly impact care received for endometrial cancer (EC) patients in Massachusetts (MA). Using 2015-2019 American Community Survey data, we implemented a Bayesian multivariate Bernoulli mixture model to identify neighborhoods with similar SDoH features in MA. Five neighborhood SDoH (NSDoH) profiles were derived and characterized: (1) advantaged non-Hispanic White; (2) disadvantaged racially/ethnically diverse, more renter-occupied housing with limited English proficiency;(3) working class, lower educational attainment; (4) racially/ethnically diverse and greater economic security and educational attainment; and (5) racially/ethnically diverse, more renter-occupied housing with limited English proficiency. Based on residential information, we assigned these profiles to EC patients in the Massachusetts Cancer Registry. We used these profile assignments as the primary exposure in a Bayesian logistic regression to estimate the odds of receiving optimal EC care, adjusting for patient-level sociodemographic and clinical characteristics. NSDoH profiles were not significantly associated with receiving optimal EC care. However, compared to patients assigned to Profile 1, patients in all other profiles had lower odds of receiving optimal care. Our findings demonstrate how a flexible model-based clustering approach can account for the interconnected and multidimensional nature of NSDoH in a practical and interpretable way. Deriving and geospatially mapping NSDoH profiles may allow for identifying areas of need and inform targeted public health interventions tailored to each neighborhood’s specific social determinants to improve healthcare delivery.
@@ 1 Background
Social determinants of health (SDoH) are defined as the social or demographic components of one’s environment that shape health outcomes and can impact an individual’s access to care, treatment decisions, and overall patient experiences [1, 2, 3]. These determinants, including socioeconomic status, experiences of racism or discrimination, and neighborhood conditions, can introduce bias in the healthcare system, potentially leading to a lack of guideline-concordant treatment [4], such as for endometrial cancer (EC), one of the most commonly occurring female cancers. Despite the existence of evidence-based treatment guidelines, disparities across the EC care continuum persist, particularly among ethnically and socioeconomically disadvantaged groups [5, 6, 7, 8, 9, 10, 11, 12, 13]. This disparate relationship serves as our motivation to better understand how SDoH patterns are considered and characterized [11, 10]. A comprehensive analysis of SDoH that transcends race and ethnicity would provide researchers and policymakers with a deeper understanding of how the social environment can influence a patient’s ability to engage with the healthcare system [9, 12].
@@
Characterization of SDoH is often conducted using non-standardized composite scores of proxy variables that correspond to different domains, such as income, education, and employment [14, 15]. At the geographic level, SDoH has been operationalized using area-based measures derived from publicly available data sources such as the United States (US) Census and the American Community Survey (ACS). Two commonly used measures are the Yost Index and the Area Deprivation Index (ADI) [16, 17, 18]. The Yost Index, typically used in national cancer surveillance, operates under the assumption that SDoH is one-dimensional. It inputs seven area-level variables (median household income, education, proportion of households below the 150% poverty line, median house value, median gross rent, unemployment rate, and working-class occupation). Summarized at the census tract or block level, these variables are clustered via a Principal Component Analysis (PCA), based on how much the respective variable explains the variation in the data[16, 19, 20]. Referred to as the primary component, the socioeconomic status (SES) index is calculated for each areal unit based on the variable loadings on that component. Areal units are then ranked based on their SES scores and divided into quintiles, with quintile 1 representing the lowest SES and quintile 5 the highest. As a one-dimensional component for SES, a loss of information results because this measure focuses on the one or two variables that may be most relevant to classifying low or high SES, as opposed to considering the combination of the different variables and how their contributions may differ from one residential area to another.
@@
The ADI, originally developed in 2002 to study general mortality disparities and hospitalizations in the US, expands the set of census-derived SDoH variables to 17. Previously, it operated under the same one-dimensional assumption as Yost, but now incorporates a three-factor model to accommodate multi-dimensionality [21, 20, 18, 17]. Highly correlated variables are first clustered together and then defined along the following three domains: financial strength, economic hardship and inequality, and educational attainment [18]. As in the Yost Index, areal units are indexed based on how high they load on the three-factor domains. Higher percentile ranks indicate greater socioeconomic disadvantage. A limitation of the ADI is that the derived domains are treated independently from one another without acknowledging the ways in which they intersect. Boscoe et al. compared the Yost and ADI indices and found that, despite being derived from the same data and heavily influenced by financial domains, they differ significantly in terms of transparency for reproducibility, distribution, and sensitivity to missing data [20]. Methodologically, these approaches share similarities in their use of principal components or factor analysis to reduce the number of highly correlated SDoH variables, but the domains derived are treated independently from one another and fall short of acknowledging the ways in which they intersect.
@@
Recently, Kolak et al. introduced a new approach that further expands the list of census-derived SDoH variables for PCA, including variables like limited English proficiency [3]. These variables were reduced to four domains: socioeconomic advantage index, limited mobility index, urban core opportunity index, and mixed immigrant cohesion and accessibility index. Additionally, they clustered census tracts by characterizing them into profiles based on similar scoring across the four SDoH components using K-means clustering. While this approach does allow for clustering based on shared responses to the different domains, the clustering is based on a single summary score. Overlooking how each component-specific score contributes to the summary score ignores the heterogeneity present in area deprivation. Two neighborhoods could potentially be clustered together but experience different patterns of resource allocation, which alter the level of impact these residents may experience in regards to their access and engagement with the healthcare system. New methodological approaches are imperative to accommodate the interconnectedness and multifactorial structure of SDoH, inclusive of factors pertinent to vulnerable populations that are at greater risk of health disparities [20, 3, 14, 22].
@@
Model-based clustering, specifically finite mixture models (FMM), are able to identify underlying patterns in heterogeneous populations from a wide set of interrelated variables. Lekkas et al. recently highlighted the methodological advantages of FMMs in neighborhood health research through the common application of Latent Class Analysis (LCA) and Latent Profile Analysis (LPA). These models are able to characterize neighborhood profiles based on a set of multivariate categorical or continuous data and examine their association with various health indicators [23]. Although LCA or LPA are well-suited to handle the multi-faceted nature of neighborhoods, typically, the number of latent patterns/profiles is unknown, and multiple model fitting and testing is required to determine the optimal number of classes. However, under a Bayesian nonparametric framework, we are able to rely on dynamic and data-driven algorithms to identify an interpretable and reproducible set of neighborhood SDoH (NSDoH) profiles with the added flexibility of incorporating prior information into parameter estimation [24]. These models have found great utility in nutrition, environmental health, and population genomics [25, 26, 27, 28, 29, 30, 31]. Yet, none have been applied to area-level exposures of SDoH.
@@
In this study, we aim to generate and characterize new NSDoH profiles to better understand their impact on EC care in Massachusetts. Given binarized aggregate proportions of area-level SDoH variables from the ACS, we will implement a Bayesian multivariate Bernoulli mixture model (MBMM) to cluster neighborhoods that share similar SDoH characteristics. Neighborhoods are established based on census tract identifiers, as defined by the American Community Survey from 2015 to 2019. Once identified, these profiles will be assigned to patients from the Massachusetts Cancer Registry, based on their residential data, and analyzed to determine if the care received was associated with their residential neighborhood.
@@
We organize the remaining sections of this paper as follows. In section 2, we describe our methodological approach, introducing the MBMM. In section 3, we apply the MBMM to our census-tract level data and examine its association with EC treatment. In section 4, we discuss the results and implications of this new approach to the field.
@@ 2.1 Data Source
We obtained census tract (neighborhood) level aggregate data from the 2015-2019 five-year American Community Survey (ACS) estimates for the state of Massachusetts (MA). The data contained a total of 1478 census tracts, which are hereafter referred to as neighborhoods, that included approximately 6.8 million persons. Based on previous research, we selected fourteen (14) variables as indicators of NSDoH [20, 3]. Given that the selected NSDoH variables were aggregate proportions and the Bernoulli distribution models binary data, we dichotomized each variable based on whether that feature fell above $(x_{ij}=1)$ or below $(x_{ij}=0)$ the state median value (Table 1). Dichotomization using the median was chosen due to the skewed distribution of several variables, which would have made other cutoffs (e.g., mean or quantile-based groupings) less interpretable or potentially biased by extreme values (Supplementary Figure 1.1). The median provides a robust, non-parametric threshold that ensures a balanced comparison between neighborhoods with relatively higher and lower exposures. Neighborhoods above the state median threshold were classified as high exposure; those below were classified as low exposure. Data was extracted from respective ACS tables via the tidycensus package implemented in the R software environment [32].
The 14 variables selected for analysis fall within four thematic domains: (1) housing conditions and resources [33, 34, 35]; (2) economic security [18, 36, 37, 38], (3) educational attainment, and (4) social and community context [3, 18]. Housing conditions and resources are described by the proportion of neighborhood households that were renter-occupied, lacked access to a vehicle, lacked complete plumbing, and experienced household crowding. Economic security is described by the proportion of households earning below the state median family income, participating in federal assistance (e.g., food stamps or SNAP - Supplemental Nutrition Assistance Program), as well as the unemployment rate, working-class status, and female head of households. Educational attainment is described by the proportion of residents with no high school diploma. Social and community context is described by the dynamic interplay between multi-ethnic communities and the social structures that shape daily life [39, 40]. We include these indicators intentionally as they can be considered proxies for racism, xenophobia, and bias, which would impact a resident’s access to resources, opportunities, and social standing [41, 42, 43]. These indicators include the proportion of households with limited English (EN) proficiency and the proportion of residents identifying as Hispanic/Latino, non-Hispanic Black, and non-Hispanic Asian. To maintain interpretability across NSDoH variables, variables were recoded to reflect indicators of greater socioeconomic disadvantage. For example, the median family income variable was recoded such that a value of $(x_{ij}=1)$ represents low income (i.e., below the state median).
@@ 2.2 Multivariate Bernoulli Mixture Model (MBMM)
Neighborhood social determinants of health (NSDoH) profiles were identified using a fully Bayesian estimation of a multivariate Bernoulli mixture model (MBMM) [44]. This model clusters a population based on shared responses to a set of binary exposures. Let $i\in\{1,...,n\}$ index an individual census tract, where $n$ is the total number of neighborhoods in Massachusetts. Let $\mathbf{x_{i}}=\{x_{i,1},..,x_{i,p}\}$ denote a vector of observed binary indicators, where $x_{i,j}=1$ denotes a high exposure to social determinant $j\in\{1,..,p\}$ in census tract $i$. Let $K$ denote the number of NSDoH profiles in the population. In practice, this number is typically not known a priori. To determine an appropriate number of profiles in the model, we overfit the model with $K_{max}$ clusters, which over-exceeds the true number of clusters. This approach, coupled with a prior that treats the number of clusters $K<K_{\text{max}}$ as an unknown parameter, allows the observed data to drive the number of nonempty clusters estimated at the end of the Bayesian sampling algorithm [45, 24, 46].The observed likelihood for the set of binary data $\mathbf{X}=\{0,1\}^{n\times p}$ under the MBMM is given by
@@ 2.2.1 Posterior Computation
Model parameters are estimated via a Metropolis-coupled Markov chain Monte Carlo ($MC^{3}$) algorithm described and implemented by Panagiotis Papastamoulis and Magnus Rattray as the R package BayesbinMix [44]. An allocation sampler is used to determine the most probable number of NSDoH profiles. This is based on the maximum a posteriori number of nonempty clusters from the posterior distribution,$K_{map}$. Once $K_{map}$ is determined, we calculate the posterior mean of each model parameter. These posterior estimates are then used to describe each NSDoH profile $k$, based on the posterior distribution of $\boldsymbol{\theta_{\cdot|k}}=(\theta_{1|k},\ldots,\theta_{p|k})$, and assigning each MA census tract to the profile with the greatest posterior probability of membership.
We assume no prior knowledge of parameter values and initialize them using noninformative priors. To ensure flexibility in capturing the complex, heterogeneous patterns of census tract-level data, we set a relatively large upper bound on the number of clusters, $K_{\text{max}}=50$, which intentionally exceeds the number of clusters we expect to identify in this data setting, which avoids underfitting.
The following noninformative priors were imposed on the model parameters to allow the data to drive estimation:
$\boldsymbol{\pi}\mid K\sim\text{Dirichlet}(\gamma_{1},\dots,\gamma_{K}),\quad\text{where }\gamma_{k}=1\;\forall\;k.$
The generated MCMC samples were postprocessed using the Equivalence Classes Representatives (ECR) algorithm to overcome label-switching identifiability issues inherent in Bayesian mixture models [47]. Further details on model fit testing and other cluster postprocessing procedures are provided in Supplemental Materials Section 1.
@@ 2.2.2 Regression Analysis
Patient-level data was obtained from the Massachusetts Cancer Registry (MCR) of the Massachusetts Department of Public Health (MDPH). A total of 2412 records were collected for women, aged 18 years or older, and diagnosed with endometrial cancer between 2015 and 2017. Cases were identified using the International Classification of Diseases for Oncology, third edition (ICD-O3) primary site (C54.1) and morphology codes for endometrial carcinoma [48]. The main outcome variable was completion of recommended treatment according to the National Comprehensive Cancer Network (NCCN) guidelines as defined in the year of treatment for each patient. These guidelines are based on tumor, grade, and stage, incorporating a combination of surgery, chemotherapy, and radiation as necessary. Adherence to NCCN treatment guidelines was treated as a dichotomous variable, where we defined optimal care as the patient receiving therapy following NCCN guidelines. Patients were assigned to a respective MBMM-derived NSDoH profile based on their residential data collected at the time of diagnosis. Associations were measured by adjusted odds ratio via a Bayesian logistic regression model, which also adjusted for patient-level clinical and sociodemographic characteristics at time of diagnosis (year of diagnosis, age, insurance status, and initial point of care facility type). Supplementary analysis was conducted to examine the association between NSDoH profiles and the type of initial care facility (academic medical center versus other facilities). More details about the guidelines and regression results are provided in Supplemental Materials Section 2.
@@ 3.1 NSDoH Profiles Results
Table 1 shows the NSDoH variables selected from the 2015-2019 ACS survey and clustered to characterize neighborhoods in MA. Neighborhoods were mostly composed of owner-occupied households and high proportions of residents with at least a high school degree. On average, across all neighborhoods, about 13% of households received federal assistance (i.e., SNAP) in the past 12 months, and 52% of the population aged 16 or older had non-white collar occupations.
@@
The model identified five NSDoH profiles, with a median profile assignment probability for neighborhoods of 0.95 (IQR: 0.18). Figure 1 illustrates the NSDoH profile patterns derived from the MBMM. To facilitate interpretation, each profile was assigned a descriptive label reflecting the predominant NSDoH characteristics. Profile 1, advantaged non-Hispanic White (NHW), represented about 32% of MA neighborhoods with very low probabilities of exposure to poor NSDoH conditions across all domains of housing, employment, education, and social context (6.5% - 37.4%). Specifically, this profile could be described as containing neighborhoods mostly comprised of owner-occupied households, and median household incomes above or equal to the state’s median. Profile 2, disadvantaged racially/ethnically diverse (BHL+; non-Hispanic Black (B) and Hispanic/Latino (HL)), more renter-occupied housing with limited EN proficiency, represented the second largest cluster, containing 25% MA neighborhoods. This profile exhibited characteristics of disadvantage and deprivation across all thematic domains. Neighborhoods in profile 2 had the highest probability of exposure to households with limited English proficiency (94.8%), female head of households (93.1%), household crowding (82.3%), renter-occupied housing (97.3%), no vehicle access (96.5%), working class occupations (98.2%), and higher proportions of all three ethnic minority groups, especially the proportion of non-Hispanic Black (81.4%) and Hispanic/Latino residents (95.2%). Profile 2 also contained neighborhoods with more residents exposed to low economic security (federal assistance participation - 99.3%; median household income below the state’s median - 99.3%). Similar trends were found in the education domain, with all reporting no HS diploma (99.4%), compared to all other NSDoH profiles. Profile 3, working class lower educational attainment, favored neighborhoods with high probability of no high school diploma (67.7%), low probabilities of exposure to poor housing conditions and resources variables, and moderate proportions of ethnic minorities (26.8-43.9%). Profile 4,racially/ethnically diverse (A+;non-Hispanic Asian (A)) and greater economic security and educational attainment, shared similar low exposures to disadvantages in the economic security and educational attainment domains with Profile 1, but differed across housing conditions and resources and social and community context. For example, Profile 4 had more neighborhoods with a higher proportion of renter-occupied housing(63.6%) and residents identifying as non-Hispanic Asians (NHA;86.7%). Profile 5, racially/ethnically diverse (ABHL+), more renter-occupied housing with limited EN proficiency, favored neighborhoods with high probabilities of exposure to crowded and renter-occupied housing, less than HS education, and ethnic minorities (84-95%). Compared to Profile 2, Profile 5 had a higher proportion of non-Hispanic Asians and less exposure to poor economic security.
Figure 1: MBMM derived NSDoH profiles. Bars indicate the estimated posterior probability of high exposure to the NSDoH variable for a neighborhood, given assignment to an NSDoH profile. SDoH variables are ordered/colored by the NSDoH thematic domain. Abbreviations: high school(HS), EN(English), Non-Hispanic (NH), Supplemental Nutrition Assistance Program (SNAP). Shorthand names for NSDoH Profiles: 1) advantaged NHW; 2) disadvantaged racially/ethnically diverse (BHL+), more renter-occupied housing with limited EN proficiency;3)working class lower educational attainment; 4) racially/ethnically diverse (A+) and greater economic security and educational attainment; 5)racially/ethnically diverse (ABHL+), more renter-occupied housing with limited EN proficiency.